OpenStudy (anonymous):

What are the values of the function y = –2x – 4 for x = 0, 1, 2, and 3?

jimthompson5910 (jim_thompson5910):

plug in x = 0 y = -2x - 4 y = -2(0) - 4 y = 0 - 4 y = -4 So y = -4 when x = 0

jimthompson5910 (jim_thompson5910):

plug in x = 1 y = -2x - 4 y = -2(1) - 4 y = -2 - 4 y = -6 So y = -6 when x = 1

jimthompson5910 (jim_thompson5910):

I'll let you do the rest

OpenStudy (anonymous):

ok thx

jimthompson5910 (jim_thompson5910):

np